OpenStudy (anonymous):

determine whether the given number is a solution to the equation. 13t + 4 = -113; -9 can I please get a step by step instruction?

OpenStudy (whpalmer4):

It couldn't be easier :-) You want to know if \(t = -9\) is a solution to the equation \[13t+4=-113\]Plug it in and find out! \[13(-9) + 4 = -113\]\[-117+4 = -113\]You tell me, is that going to be true?

OpenStudy (whpalmer4):

Let's try a different one. What if we thought \(t = 9\) was also a solution? \[13(9) + 4 = -113\]\[117+4 = -113\]\[121=-113\]Oops, that's not a solution!

OpenStudy (anonymous):

wow thanks so much I got it!!

OpenStudy (whpalmer4):

Great! it's an invaluable technique. Do some work, get an answer, try it, see if it appears to answer the question correctly! Helpful tip: before you do the check, reread the problem again to make sure you've really answered the right question, too. Easy to get tangled up in a "word problem" and end up finding the answer to some related question, rather than the one you need to answer.